# On Finding All Connected Maximum-Sized Common Subgraphs in Multiple Labeled Graphs

### v1

@@ 4 Results
To test the effect of ordering the input graphs together with “refining the type-A edges” as outlined in the last part of paragraph Techniques for finding type-A connected cliques, we we created 5000 problem instances, each containing 5 molecular graphs with 35 atoms (excluding hydrogen) randomly chosen from the ChEMBL22 database [19]. All experiments were performed on a machine with 32 GB of LPDDR5X RAM, an Intel Core Ultra 9 processor and 1TB NVMe SSD, running the Fedora Workstation 41 operating system. To determine an ordering of the input graphs we used the four similarity measures VH, WL, NSPD and minmax and the method outlined in the paragraph Ordering of Input Graphs. For each ordering, we determined the maximal type-A connected cliques with and without the refinement of type-A edges. Each of the eight different configurations have been applied in the 5000 test instances and the run-time in seconds has been recorded. The resulting box plots are shown in Figure 2. We emphasize that Figure 2 does not include box plots for randomly ordered graphs because, in many cases, the runtime for computing the maximal type-A connected cliques exceeded one hour ($3.6\times 10^{3}$ seconds)—if they finished at all. Furthermore, we highlight that the ordering used in the approach established in [7] is based on the size of the molecular graphs. However, since all test instances contain 35 vertices, this ordering is expected to be essentially random. As a result, its runtime can also be assumed to exceeded one hour—if it finished at all. In contrast, the mean runtime for graphs ordered by similarity ranges from $0.25$ to $1.99$ seconds, highlighting the significant impact of the ordering methods. Moreover, it can be observed that the runtime improves in all cases when applying the refinement of the type-A edges. The minmax ordering with the applied refinement step achieves the lowest mean runtime. Furthermore, even without the refinement step, minmax ordering still results in the lowest mean runtime compared to the other ordering methods, regardless of whether refinement is applied.
@@ Appendix 0.D Related works
A further complication in the MCS domain is the distinction between two related formulations. Many methods target the maximal common induced subgraph (MCIS), defined as the largest vertex-induced subgraph common to the compared molecules, while other methods focus on the maximal common edge subgraph (MCES), which maximizes the number of shared edges [30, 31]. This duality has led to a generalized classification framework for MCS algorithms that not only categorizes them as exact or approximate but also distinguishes whether the resulting subgraph is required to be connected or may be disconnected [9, 32, 33, 34].

### v2

@@ 4 Results
To test the effect of ordering the input graphs together with “removal of certain type-0 edges” for finding MECS as outlined in the last part of paragraph Techniques for finding type-A connected cliques, we created 5000 problem instances, each containing 5 molecular graphs with 35 atoms (excluding hydrogen) randomly chosen from the ChEMBL22 database [19]. All experiments were performed on a machine with 32 GB of LPDDR5X RAM, an Intel Core Ultra 9 processor and 1TB NVMe SSD, running the Fedora Workstation 41 operating system. To determine an ordering of the input graphs we used the four similarity measures VH, WL, NSPD and minmax and the method outlined in the paragraph Ordering of Input Graphs. For each ordering, we determined the maximal type-A connected cliques with and without the removal of certain type-0 edges. Each of the eight different configurations have been applied in the 5000 test instances and the run-time in seconds has been recorded. The resulting box plots are shown in Figure 2. We emphasize that Figure 2 does not include box plots for randomly ordered graphs because, in many cases, the runtime for computing the maximal type-A connected cliques exceeded one hour ($3.6\times 10^{3}$ seconds)—if they finished at all. Furthermore, we highlight that the ordering used in the approach established in [8] is based on the size of the molecular graphs. However, since all test instances contain 35 vertices, this ordering is expected to be essentially random. As a result, its runtime can also be assumed to exceeded one hour—if it finished at all. In contrast, the mean runtime for graphs ordered by similarity ranges from $0.25$ to $1.99$ seconds, highlighting the significant impact of the ordering methods. Moreover, it can be observed that the runtime improves in all cases when applying the removal of the specified type-0 edges. The minmax ordering with the applied refinement step achieves the lowest mean runtime. Furthermore, even without the refinement step, minmax ordering still results in the lowest mean runtime compared to the other ordering methods, regardless of whether refinement is applied.
@@ Appendix 0.D Related works
A further complication in the MCS domain is the distinction between two related formulations. Many methods target the maximal common induced subgraph (MCIS), defined as the largest vertex-induced subgraph common to the compared molecules, while other methods focus on the maximal common edge subgraph (MCES), which maximizes the number of shared edges [30, 31]. This duality has led to a generalized classification framework for MCS algorithms that not only categorizes them as exact or approximate but also distinguishes whether the resulting subgraph must be connected or may be disconnected [32, 33, 34, 35].
@@
Acknowledgements
This work was funded by the Novo Nordisk Foundation as part of the project MATOMIC (Mathematical Modelling for Microbial Community Induced Metabolic Diseases, 0066551) and from the European Unions Horizon Europe Doctoral Network programme under the Marie-Sklodowska-Curie grant agreement No 101072930 (TACsy – Training Alliance for Computational systems chemistry).